OpenStudy (anonymous):

The line passing through the points (-1, 6) and (4, b) has gradient -2. Find the value of b.

OpenStudy (anonymous):

using the point-slope equation: y - y1 = m(x - x1) b - 6 = -2(4 -1) b - 6 = -6 b = 0

OpenStudy (anonymous):

welcome to OS ! :)

OpenStudy (skullpatrol):